OpenStudy (anonymous):

A town's public works department has ordered sand to spread on roadways during the upcoming winter snows. The delivered sand forms a right cone 36 feet high and a base diameter of 54 feet. How many cubic yards of sand has the department received?

OpenStudy (anonymous):

A cone is 1/3 of a cylinder the formula is V = (1/3) pi * r^2 * h so we have V = (1/3) * pi * 27^2 * 36 = 27482.65253 cubic feet

OpenStudy (anonymous):

However, 27 cubic feet = 1 cubic yard so 27482.65253/27 = 1017.87602 cubic yards

OpenStudy (anonymous):

oh thanks can you help me with quite a few more please

OpenStudy (anonymous):

A contractor has installed a silt fence around an area that is semi-circular and level to prevent soil from the construction site entering nearby streams. The diameter of the semi-circle is 1200 feet. How many linear feet of fence does the contractor need to use to enclose the area?

OpenStudy (anonymous):

Does the fence connect the two ends of the semicircles?

OpenStudy (anonymous):

semicircle*

OpenStudy (anonymous):

Anyways, you'd find the circumference of a circle (2*pi*r). Semi-circle is half of a circle so you can use (pi*r). The radius is 600 ft, so for the semi-circular part of the fence you'd need approximately 1884.96 linear ft of fence. If you connect the two ends of the semi-circle, add another 1200 linear feet so you'd have 3084.96 linear ft.

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

how many acres would be enclosed by the fence?

OpenStudy (anonymous):

You find the area of the semicircle then. (pi * r^2)/2. Then convert to acres. 1 acre = 43,560ft^2

OpenStudy (anonymous):

oh wow thanks once again. i have about 3 more questions, do you have the time?

OpenStudy (anonymous):

Just post them, I'm sure someone will answer.